Factorise;
axy-ay+ax-a
I think I've got the correct answer;
(ay+a)(x-1)
but I kind of just did trial and error. I was wondering if there's a faster or more straight forward way of getting there?
Thanks SO much again (you guys are AWESOME!)
manda

Found 2 solutions by edjones, unlockmath:
Answer by edjones(8007)

You almost got it.
axy-ay+ax-a
=ay(x-1)+a(x-1) factoring by grouping.
=(ay+a)(x-1)
=a(y+1)(x-1)
.
Ed

Answer by unlockmath(1688)

Hello Manda,
You are correct. Good going! We can go one more step. (See below)
Since I didn't see how you did it I couldn't say if your way was fastest. Here's the way I'd do it. I see in the first two terms a common thread of ay, so we could pull that out like:
ay(x-1)
Now notice in the last two terms the common thread is a, so we can rewrite that:
a(x-1) Notice that these two have (x-1) therefore we can write this as:
(ay+a)(x-1) We could go one more step and write it as:
a(y+1)(x-1) Multiply all that out and we should arrive at the original expression.
RJ Toftness
www.math-unlock.com